OpenStudy (anonymous):

A=1/2h(B+b) solve for h A=85,B=10, b=10

OpenStudy (anonymous):

Well plug everything i.

OpenStudy (anonymous):

in*

OpenStudy (anonymous):

85=1/2h(10+10)

OpenStudy (anonymous):

85=1/2h(20)

OpenStudy (anonymous):

85=h10

OpenStudy (anonymous):

h=85/10

OpenStudy (anonymous):

h=8.5